OpenStudy (anonymous):

Mr. Crosson selects two students from his class to present their reports on Spain. He places the names of each student in a box and picks one at a time. The class has 8 boys and 10 girls. What is the probability that Mr. Crosson will pick two boys in a row if he does not put the first name back in the box? Round your answer to the nearest percent.

OpenStudy (kropot72):

The total number of names is 8 + 10. On the first draw the probability of a boy is \[P(b _{1})=\frac{8}{8+10}\] On the second draw there are 7 boys names and a total of 7 + 10 names. the probability of a boy is \[P(b _{2})=\frac{7}{7+10}\] The two draws are independent events so the probability of two boys is \[P(2\ boys)=\frac{8}{8+10}\times \frac{7}{7+10}=you\ can\ calculate\]